# Symbolic Regression with a Learned Concept Library

### v1

@@ Abstract
We present a novel method for symbolic regression (SR), the task of searching for compact programmatic hypotheses that best explain a dataset. The problem is commonly solved using genetic algorithms; we show that we can enhance such methods by inducing a library of abstract textual concepts. Our algorithm, called LaSR, uses zero-shot queries to a large language model (LLM) to discover and evolve concepts occurring in known high-performing hypotheses. We discover new hypotheses using a mix of standard evolutionary steps and LLM-guided steps (obtained through zero-shot LLM queries) conditioned on discovered concepts. Once discovered, hypotheses are used in a new round of concept abstraction and evolution. We validate LaSR on the Feynman equations, a popular SR benchmark, as well as a set of synthetic tasks. On these benchmarks, LaSR substantially outperforms a variety of state-of-the-art SR approaches based on deep learning and evolutionary algorithms. Moreover, we show that LaSR can be used to discover a novel and powerful scaling law for LLMs.
@@ 4 Experiments
LaSR’s main focus is to serve as a practical toolkit for scientists. Therefore, our evaluation primarily targets slightly noisy environments, using exact solution rate to gauge performance rather than statistical similarity measures like correlation $R^{2}$, which are less relevant to scientific discovery applications.
@@ 4.1 Comparison against baselines in the Feynman Equation Dataset
Dataset: The Feynman Equations dataset is a well established benchmark for Symbolic Regression algorithms [47]. This dataset consists of 100 physics equations extracted from the Feynman lectures on Physics. We compare against performance reported on SRBench [25]: a continuously updated public benchmark of SR methods on many datasets. Specifically, we compare against GPlearn, AFP, AFP-FE, DSR, uDSR, PySR, and AI Feynman [40, 45, 47, 27, 36]. Within this subset PySR represents an ablation of our model without the LLM genetic operations and the concept evolution (Section 3). We evaluate on a slightly noisy version of this dataset in order to simulate experimental errors common in scientific discovery domains. Details about this dataset are presented in Section A.4.1.
@@ 4.3 Ablation Experiments
We conduct ablations on the use of Concept Evolution, Concept Crossover, variable names, and user hints. Figure 3 shows how these ablations affect performance over 40 iterations. We designate an equation as “solved” if, after $N$ iterations, the MSE of our predicted equation is less than $10^{-11}$. This metric differs from ’Exact Solved’ as defined in the prior experiments: an equation can be ’exactly solved’ yet have an MSE higher than $10^{-11}$ due to the noise floor in the target variables, and an equation can have low loss but not be an exact match. We observe from the results that: (1) Removing variable names results in a substantial performance drop, as we lose semantic meaning provided by variables (for instance, observing $\theta$ could suggest employing trigonometric functions on $\theta$). (2) Learning a concept library enables faster convergence to solutions. Without the concept library, task convergence is slower, and widens under higher concept guidance conditions ($p>0.1\%$).
@@ 4.4 Qualitative Analysis and User Hints
Figure 3: Evaluation results for ablations/extensions of LaSR. (Left): We ablate three components of LaSR: Concept Evolution, Concept Crossover, and variable names and evaluate their MSE solve rate performance on the Feynman dataset over 40 iterations. We find that each component contributes to accelerating search at different stages in the search process. (Right): We extend LaSR by providing an initial concept library $\mathcal{C}_{0}$ in the form of user provided hints. We find that natural language hints significantly increases the speed of solving equations.
@@ 4.6 Using LaSR to discover LLM Scaling Laws
Qualitative Evaluation: Equation 4 describes an empirical relationship between training hyperparameters (training steps) and inference hyperparameters (number of shots). It asserts that increasing the number of shots exponentially increases the model’s performance for low-resource models, while having diminishing gains as the number of training steps of the model increase. This observation is consistent with work in scaling test-time compute [44]. However, if the number of shots is an odd number, the exponentially increasing trend reflects over the line $\texttt{score}=E$ and turns into an exponentially decreasing trend. A manual inspection of the training data reveals the majority of samples had an even number of shots (2, 4, …) which causes this generalization error.
@@ 5 Related Work
The field SR started in the 1970s [17, 28] and has recently become a prominent approach to AI-for-science [32, 33, 38]. Two algorithmic themes here are:
@@
Parametric Algorithms: Recent work in SR and program synthesis has often used neural networks to accelerate search [41, 38, 36, 27, 33, 12]. The interplay between the neural and the symbolic components in these works can be abstracted into two categories: (1) leveraging LLMs to induce program scaffolds [33, 38], and (2) learning a neural policy to accelerate search [36, 27, 41, 12]. We highlight two methods from the first category: Funsearch [38] and LLM-SR [43]. Funsearch [38] uses a pretrained LLM to implement a mutation operator on a database of executable programs under a fixed specification to find super-optimized programs in extremal combinatorics. LaSR is a generalization of FunSearch: while FunSearch conditions program generation on a static “specification” (analogous to our concept library), we discover the concept library in the course of the algorithm. As for LLM-SR [43], it leverages a pretrained LLM for generating program sketches [34]. The sketch parameters are optimized and cached in a database which is in turn used to generate new sketches. Our work is an orthogonal direction of improvement. It is technically possible to “plug” the LLM-SR framework into LaSR and use our generated ideas to guide the lower-level search component.
The second category includes methods like DSR [36], which, just like LaSR, frames SR as a sequence modeling problem. However, the search in LaSR leverages a learned concept library and the language and code biases in LLMs, instead of relying on amortization alone.
@@
Recent work in program synthesis models program generation as a sequence prediction problem. Under this paradigm, the DSL and the input-output specification is serialized in the prompt and a code-generation foundation model [30, 7, 4] is leveraged to autoregressively generate candidate programs. This approach has been used to successfully synthesize programs in many areas including spreadsheet formula prediction [12, 8], competitive programming [31], and visual programming [46, 20]. LaSR is similar to work in this area in that the LLM Mutate, LLM Crossover, and LLM Initialization functions all follow the sequence prediction paradigm to synthesize mathematical equations, relying on guidance from the concept library.
@@ A.2 LLM Prompts
Figure 5: LLM Concept Abstraction prompt with an example output. The LLM Concept Crossover function follows a similar structure with a modified task description for crossover on concepts.
@@ A.3.1 Compute Usage
We run all experiments on a server node with 8xA100 GPUs with 80 GB of VRAM each. However, our experiments can be reproduced with a GPU with 16 GB of VRAM. We were even able to run LaSR on a laptop utilizing a quantized model hosted locally  [^2] TheBloke/Mistral-7B-Instruct-v0.2-GGUF using llama.cpp. Moreover, certain models are hosted on external servers (such as gpt-3-turbo-0125) which allows running LaSR on machines without GPUs. For this project, we chose to run llama3-8b using vLLM [24]. However, our framework is compatible with any LLM inference framework that allows hosting an OpenAI compliant RESTful server. For reference, each iteration makes around $60,000$ calls. Each call to the LLM is just under $1000$ tokens. This gives an upper bound on total compute of $60,000,000$ tokens per iteration if $p=100\%$. Hence, running our model at $p=1\%$ for 40 iterations would result in just under 25M tokens for each equation.
@@ A.3.3 Hyperparameters
Figure 7 showcases the hyperparameters used for all our experiments. Wherever possible, we use the default PySR parameters. Additionally, LaSR introduces three new hyperparameters: (1) % of LLM calls, (2) List of user hints, and (3) a dictionary of parameters pertaining to backend LLM communication. Following other methods in SRBench, we utilize only a subset of the necessary operators for solving the Feynman equations, excluding special operators like $\arcsin$ and $\arctan$. These operators are seldom required, and removing them speeds up the search process. We generally set the number of iterations to $40$. However, certain experiments may demand more or less iterations.
@@ A.4.2 Synthetic Dataset
A.5 Using LaSR to find an LLM Scaling Law
@@ Results:
A.6 Metrics for Cascading experiment
For the cascading experiment, we aim to evaluate the progression of different configuration towards solving equations. The quantitative metric used in the SRBench comparision experiment, Exact Solve, does not allow for such fine-grained analysis. Therefore, we categorize the synthesized equations into four buckets: Exact Solve, Almost Solve, Close, and Not Close. Exact Solve is quantitatively evaluated using a symbolic match. An equation is tagged as ’Almost Solve’ if the dataset loss is small, but the generated equation has an extra term or lacks one term. A Close equation captures the general structure of the solution (such as a square root nested in an exponential) but not more than that, and Not Close includes all equations that are far from the solution.
A.7 Further Qualitative Analysis
@@
A.8 Subset of equations discovered by LaSR

### v2

@@ Abstract
We present a novel method for symbolic regression (SR), the task of searching for compact programmatic hypotheses that best explain a dataset. The problem is commonly solved using genetic algorithms; we show that we can enhance such methods by inducing a library of abstract textual concepts. Our algorithm, called LaSR, uses zero-shot queries to a large language model (LLM) to discover and evolve abstract concepts occurring in known high-performing hypotheses. We discover new hypotheses using a mix of standard evolutionary steps and LLM-guided steps (obtained through zero-shot LLM queries) conditioned on discovered concepts. Once discovered, hypotheses are used in a new round of concept abstraction and evolution. We validate LaSR on the Feynman equations, a popular SR benchmark, as well as a set of synthetic tasks. On these benchmarks, LaSR substantially outperforms a variety of state-of-the-art SR approaches based on deep learning and evolutionary algorithms. Moreover, we show that LaSR can be used to discover a new and powerful scaling law for LLMs.
@@ 4 Experiments
LaSR’s main focus is to serve as a practical toolkit for scientists. Therefore, our evaluation primarily targets slightly noisy environments, using exact solution rate to gauge performance rather than statistical similarity measures like correlation $R^{2}$, which are less relevant to scientific discovery applications. Additional experiments characterizing the practical behavior of LaSR are in Appendix A.5. Information regarding compute usage is in Appendix A.3.1.
@@ 4.1 Comparison against baselines in the Feynman Equation Dataset
Dataset: The Feynman Equations dataset is a well established benchmark for Symbolic Regression algorithms [49]. This dataset consists of 100 physics equations extracted from the Feynman lectures on Physics. We compare against performance reported on SRBench [26]: a continuously updated public benchmark of SR methods on many datasets. Specifically, we compare against GPlearn, AFP, AFP-FE, DSR, uDSR, PySR, and AI Feynman [42, 47, 49, 28, 38]. Within this subset PySR represents an ablation of our model without the LLM genetic operations and the concept evolution (Section 3). We evaluate on a slightly noisy version of this dataset in order to simulate experimental errors common in scientific discovery domains. More details are presented in Appendix A.4.1.
@@ 4.3 Ablation Experiments
We conduct ablations on the use of Concept Evolution (skip phase three from Figure 1), Concept Library (skip phase two and three), variable names, and user hints. Figure 3 shows how these ablations affect performance over 40 iterations. We designate an equation as “solved” if, after $N$ iterations, the MSE of our predicted equation is less than $10^{-11}$. This metric differs from ’Exact Solved’ as defined in the prior experiments: an equation can be ’exactly solved’ yet have an MSE higher than $10^{-11}$ due to the noise floor in the target variables, and an equation can have low loss but not be an exact match. We observe from the results that: (1) Removing variable names results in a substantial performance drop, as we lose semantic meaning provided by variables (for instance, observing $\theta$ could suggest employing trigonometric functions on $\theta$). (2) Learning a concept library enables faster convergence to solutions. Without the concept library, task convergence is slower, and widens under higher concept guidance conditions ($p>0.1\%$).
@@ 4.4 Qualitative Analysis and User Hints
Figure 3: Evaluation results for ablations/extensions of LaSR. (Left): We ablate three components of LaSR: Concept Evolution, Concept Library, and variable names and evaluate their MSE solve rate performance on the Feynman dataset over 40 iterations. We find that each component contributes to accelerating search at different stages in the search process. (Right): We extend LaSR by providing an initial concept library $\mathcal{C}_{0}$ in the form of user provided hints. We find that natural language hints significantly increases the speed of solving equations.
@@ 4.6 Using LaSR to discover LLM Scaling Laws
Qualitative Evaluation: Equation 4 describes an empirical relationship between training hyperparameters (training steps) and inference hyperparameters (number of shots). It asserts that increasing the number of shots exponentially increases the model’s performance for low-resource models, while having diminishing gains as the number of training steps of the model increase. This observation is consistent with work in scaling test-time compute [46]. However, if the number of shots is an even number, the exponentially increasing trend reflects over the line $\texttt{score}=E$ and turns into an exponentially decreasing trend. A manual inspection of the training data reveals the majority of samples had an odd number of shots (1, 3, …) which causes this generalization error.
@@ 5 Related Work
The field of SR started in the 1970s [18, 29] and has recently become a prominent approach to AI-for-science [33, 34, 40]. Two algorithmic themes here are:
@@
Parametric Algorithms: Recent work in SR and program synthesis has often used neural networks to accelerate search [43, 40, 38, 28, 34, 13, 35]. The interplay between the neural and the symbolic components in these works can be abstracted into two categories: (1) leveraging LLMs to induce program scaffolds [34, 40, 35], and (2) learning a neural policy to accelerate search [38, 28, 43, 13]. We highlight two methods from the first category: Funsearch [40] and LLM-SR [45]. Funsearch [40] uses a pretrained LLM to implement a mutation operator on a database of executable programs under a fixed specification to find super-optimized programs in extremal combinatorics. LaSR is a generalization of FunSearch: while FunSearch conditions program generation on a static “specification” (analogous to our concept library), we discover the concept library in the course of the algorithm. We do not compare against FunSearch due to resource constraints. As for LLM-SR [45], it leverages a pretrained LLM for generating program sketches [36]. The sketch parameters are optimized and cached in a database, which is in turn used to generate new sketches. Our work is an orthogonal direction of improvement. It is technically possible to “plug” the LLM-SR framework (or other LLM-based search algorithms [35]) into LaSR and use our generated concepts to guide the lower-level search component.
The second category includes methods like DSR [38], which, just like LaSR, frame SR as a sequence modeling problem. However, the search in LaSR leverages a learned concept library and the language and code biases in LLMs, instead of relying on amortization alone.
@@
Recent work in program synthesis models program generation as a sequence prediction problem. Under this paradigm, the DSL and the input-output specification is serialized in the prompt and a code-generation foundation model [31, 7, 4] is leveraged to autoregressively generate candidate programs. This approach has been impactful in many areas including spreadsheet formula prediction [13, 8], competitive programming [32], and visual programming [48, 21, 9]. LaSR is similar to work in this area in that the LLM Mutate, LLM Crossover, and LLM Initialization functions all follow the sequence prediction paradigm to synthesize mathematical equations, relying on guidance from the concept library.
@@ A.2 LLM Prompts
Figure 5: LLM Concept Abstraction prompt with an example output. The LLM Concept Crossover function follows a similar structure, with a modified task description for crossover on concepts.
@@ A.3.1 Compute Usage
We run all experiments on a server node with 8xA100 GPUs with 80 GB of VRAM each. However, our experiments can be reproduced with a GPU with 16 GB of VRAM. We were even able to run LaSR on a laptop utilizing a quantized model hosted locally  [^2] TheBloke/Mistral-7B-Instruct-v0.2-GGUF using llama.cpp. Moreover, certain models are hosted on external servers (such as gpt-3-turbo-0125) which allows running LaSR on machines without GPUs. For this project, we chose to run llama3-8b using vLLM [25]. However, our framework is compatible with any LLM inference framework that allows hosting an OpenAI compliant RESTful server. For reference, each iteration makes around $60,000$ calls. Each call to the LLM is just under $1000$ tokens. This gives an upper bound on total compute of $60,000,000$ tokens per iteration if $p=100\%$. Hence, running our model at $p=1\%$ for 40 iterations would result in just under 25M tokens for each equation (around 4 hours on a single A100).
@@ A.3.3 Hyperparameters
xFigure 7 showcases the hyperparameters used for all our experiments. Wherever possible, we use the default PySR parameters. Additionally, LaSR introduces three new hyperparameters: (1) % of LLM calls, (2) List of user hints, and (3) a dictionary of parameters pertaining to backend LLM communication. Following other methods in SRBench, we utilize only a subset of the necessary operators for solving the Feynman equations, excluding special operators like $\arcsin$ and $\arctan$. These operators are seldom required, and removing them speeds up the search process. We generally set the number of iterations to $40$. However, certain experiments may demand more or less iterations.
@@ A.4.2 Synthetic Dataset
A.5 Additional Experiments
This section highlights additional experiments on various benchmarks to further characterize LaSR’s performance.
A.5.1 Asymmetric comparison with PySR
A common concern with evaluating genetic optimization algorithms w.r.t number of iterations is that the ’hard stop’ after a certain number of iterations might yield populations that haven’t fully converged to their optimal values.
To account for this discrepancy in performance, we conduct an asymmetric comparison with PySR on the Feynman Equations dataset. Specifically, we allow PySR to run uninterrupted for 10 hours per equation, with the maximum number of iterations set to $10^{6}$. We compare the results with that of LaSR run interrupted for 40 iterations per equation (hence the asymmetric comparison). The results are detailed in Table 5.
Overall, we find that, despite running PySR substantially longer than LaSR, LaSR still substantially outperforms PySR. This is because PySR, like other evolutionary algorithms, is susceptible to falling in local minima and converges to this local minima extremely fast. This is why supplementing “local” search strategies with LLM guidance is useful.
Metric PySR ($10^{6}$ Iterations, 10 hour timeout) LaSR ($40$ iterations)
Exact Solve $59+3/100$ 72/100
Table 5: An asymmetric comparison of PySR and LaSR on the Feynman equations dataset (A.5.1). We run PySR for 10 hours per equation (thresholded to $10^{6}$ iterations) and compare the exact solve rate with that of LaSR run for 40 iterations. We find that PySR is able to discover three more equations, but LaSR still substantially outperforms PySR.
A.5.2 Qualitative comparision of Synthetic Dataset equations
In Table 3, we reported $R^{2}$ test set performance of LaSR and PySR on the synthetic dataset. In this section, we qualitatively compare the equations discovered by LaSR and PySR on the procedurally generated dataset of equations. None of the algorithms recover the ground truth form, but we find that LaSR’s equations fit much better to the ground truth data than PySR’s equations. These equations are presented in Table 3.
GT Equation Predicted Equation $\exp\left(\frac{C-\log(y_{2})}{C-y_{4}}+\left(\sqrt{y_{1}+y_{4}}+\sqrt{y_{2}}\right)\right)$ $\left(\frac{(y_{4}+C+y_{1})}{C}\cdot(y_{2}+C)\right)^{C}+y_{1}$ $\frac{(y_{1}+C)+\left(\sqrt{y_{3}}\cdot\left(y_{1}\cdot\left(C-y_{3}\right)\cdot y_{3}\right)\right)}{\cos(\cos(y_{1}))}$ $y_{3}\cdot\frac{C-\left(y_{1}\cdot\left(y_{3}-\left(\cos\left((y_{1}+C)\cdot C\right)\cdot C\right)+C\right)\right)}{C}-C$ $\sqrt{\exp(y_{1})\cdot y_{1}}\cdot\left(\cos(y_{1})+2y_{1}\right)$ $\left(y_{1}^{C}+C\right)\cdot\left(\sqrt{\exp(y_{1})}-C\right)+\left(C-\left(\frac{\left(y_{1}^{2}\cdot\exp(\cos(y_{1}))\right)}{C}+y_{4}-C\right)\right)$ $\exp\left(\log(y_{4}+y_{2}\cdot y_{1})\right)\cdot\left(\cos(y_{2})-\sqrt{y_{1}}+\frac{y_{3}}{C}-C\right)$ $y_{1}\cdot\left((y_{2}\cdot\cos(y_{2}))-y_{4}-(y_{3}+1)\cdot y_{2}\right)-1$ $(y_{3}+y_{2})\cdot\left(\left((y_{3}+C)\cdot y_{2}+y_{1}+y_{3}\right)-C\right)$ $\left((y_{2}+y_{3})\cdot\left(y_{3}+\left(y_{2}\cdot\left(y_{3}+1\right)\right)+1+y_{1}\right)-C+y_{3}\right)$ $\left(\sqrt{y_{2}}+\exp\left(C+\sqrt{y_{2}}\right)\right)\cdot\left((y_{4}\cdot y_{3})\cdot\log(y_{3})\right)$ $\frac{\left(\left(\sqrt{y_{2}^{C}}-C\right)\cdot(y_{4}\cdot C\cdot(y_{3}-C))-\sin\left(y_{2}^{C}+C\right)\right)\cdot\log(y_{3})}{C}+C$ $\frac{(C\cdot y_{3})\cdot\exp(\sqrt{y_{2}}-y_{1})}{C/(y_{2}+y_{1})}$ $\left(y_{2}\cdot\frac{C-\sqrt{y_{2}^{C}}}{\left(\sqrt{y_{1}^{C}}+C\right)^{y_{1}}}\right)\cdot y_{3}-C$ $\frac{\left(y_{2}+C\right)\cdot\sqrt{y_{1}}}{C}\cdot\sqrt{\frac{\exp(y_{1})}{y_{1}}}$ $\frac{C-y_{2}}{\sin\left(\frac{C}{y_{1}+C}\right)^{C}}-\cos(C-y_{1})$ $y_{1}\cdot\left(\frac{y_{1}^{2}}{\frac{C}{\cos\left(C-y_{4}-C\right)}}\right)/\exp(C\cdot y_{5})$ $\left(\left(y_{1}-C\right)\cdot\left(\sin(y_{4}+C)\cdot y_{1}\right)\cdot(y_{5}+y_{1}-C)\right)\cdot C-C$ $\exp(\log(y_{3})\cdot\sqrt{y_{2}})-\exp\left(\sqrt{(y_{3}+y_{4})+(y_{1}\cdot y_{2})}\cdot C\right)$ $\left(\frac{y_{3}\cdot(y_{2}/C)^{C}+C-y_{4}-y_{2}}{-y_{2}}\right)$
Table 6: Qualitative evaluation of LaSR on the synthetic equations dataset (A.5.2). We attempt to recover the ground truth equation from a slightly noisy dataset generated from the ground truth equations. None of the algorithms we tested (including LaSR) are able to recover the ground truth equations, underscoring the challenge of exact symbolic match on this dataset. A study on the $R^{2}$ performance is presented in Table 3.
A.5.3 Stochasticity of LaSR and PySR
A common concern with using LLM’s for scientific discovery is that the equations could be obtained due to dataset memorization rather than pure reasoning and learning on the equations form. To explore this further, we run LaSR with the same hyperparameters twice (with different seeds). If the LLM purely relies on memorization, we expect the model to find the same equation form in both experiments. We present results in Table 7.
Discovered Equation (LaSR - Llama 3b 1%; same params) Loss Complexity
$h=\frac{2.8841133}{0.000219-\left(r\cdot\left(\frac{-36.240696-(\sin(r)\cdot 0.002057)}{P}\right)\cdot r\right)}$ $2\times 10^{-6}$ $16$
$h=\frac{P}{r\cdot(r\cdot 17.252695-0.001495)}\cdot 1.372847$ $2\times 10^{-6}$ $11$
Table 7: Performance of LaSR on successive runs with the same hyperparameters on Equation 53 in the Feynman equations dataset (A.5.3). This equation defines the heat flow from a point source (an inverse square law with ground truth formulation: $h=\frac{P}{4\pi R^{2}}$). We evaluate LaSR by running it twice with identical hyperparameters but different random seeds to assess whether the model has memorized the equation’s form. In both runs, LaSR consistently discovers high-performing solutions. Moreover, the resulting functional forms show significant variation between runs, supporting the hypothesis that LaSR’s success is not rooted in memorization.
Overall, both equations fit well to the underlying dataset and reduce to the ground truth. Yet, despite using the same hyperparameters, the functional forms exhibit sharp differences. This further reinforces our hypothesis that LaSR’s performance is not simply the result of regurgitated memorized responses.
A.6 Using LaSR to find an LLM Scaling Law
@@ Results:
A.7 Metrics for Cascading experiment
For the cascading experiment, we aim to evaluate the progression of different configuration towards solving equations. The quantitative metric used in the SRBench comparison experiment, Exact Solve, does not allow for such fine-grained analysis. Therefore, we categorize the synthesized equations into four buckets: Exact Solve, Almost Solve, Close, and Not Close. Exact Solve is quantitatively evaluated using a symbolic match. An equation is tagged as ’Almost Solve’ if the dataset loss is small, but the generated equation has an extra term or lacks one term. A Close equation captures the general structure of the solution (such as a square root nested in an exponential) but not more than that, and Not Close includes all equations that are far from the solution.
A.8 Further Qualitative Analysis
@@
A.9 Subset of equations discovered by LaSR